OpenStudy (anonymous):

There were 170 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $490. How many Juniors bought tickets? 70 80 90 100

OpenStudy (anonymous):

Let x be the number of Juniors then there are (170-x)number of Senior.... now, 2x+ 3.50(170-x) = 490

OpenStudy (anonymous):

Solve that get the value of x

OpenStudy (anonymous):

can u do tht for mr cause i m doing this for my cousin who is taking an online course for algebra 1 @sauravshakya

OpenStudy (anonymous):

*me

OpenStudy (anonymous):

Am I allowed to do that?????/

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

2x+ 3.50(170-x) = 490 2x+595-3.5x = 490 -1.5x=-105 x=70

OpenStudy (anonymous):

So, what do u think is the answer?

OpenStudy (anonymous):

x= 70

OpenStudy (anonymous):

r u indian?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

nope......

OpenStudy (anonymous):

Do i look like indian?

OpenStudy (anonymous):

no ur name @sauravshakya

OpenStudy (anonymous):

Oh........ I am from NEPAL

OpenStudy (anonymous):

oh kk